OpenStudy (kirbykirby):

Subtracting 2 random variables: Let \(X\)~\(N(0,1)\) and \(Y\)~\(N(0,1)\) be independent random variables. Show that \(X-Y\)~\(N(0,2)\).

OpenStudy (kirbykirby):

I did: Let \(X=X_1\) and let \(Y=X_2\) Then, there is a theorem stating that if \(X_i,i=1,2,...n\) are independent \(N(\mu_i,\sigma_i ^2)\) r.v.'s and \(a_i,i=1,2,...,n\) are constants, then \(\sum_{i=1}^{n}a_i X_i\)~\(N(\sum_{i=1}^{n}a_i \mu_i\,,\sum_{i=1}^{n}a_i ^2 \sigma_i ^2\)). So, \(X-Y=X_1-X_2\)~\(N((1)(0)+(-1)(0),(1)^2(1)^2+(-1)^2(1)^2)=N(0,2)\) But I find it odd because I know when we add r.v.'s,(like if we had X+Y.... can we not say X=Y because they have the same distribution, so X+X=2X? It's just odd for X-Y because that would yield X-X= 0 ??